OpenStudy (anonymous):

What is the approximate distance from (4, 6) to (-2,1)?

OpenStudy (anonymous):

\[\sqrt{(x2-x1)^2 + (y2-y1)^2)}\]

hartnn (hartnn):

the formula for distance between two points is given by ^^

OpenStudy (anonymous):

\[\sqrt{61}\]

OpenStudy (anonymous):

its eiather 5 9 6 or 8 units

hartnn (hartnn):

sqrt(36+25)=sqrt(61) nopes ,check your points or options.

OpenStudy (anonymous):

it depends how their asking the question as well.

OpenStudy (anonymous):

What is the approximate distance from (4, 6) to (-2,1)? 5 units 9 units 6 units 8 units

OpenStudy (anonymous):

find the square root of 61 then round it prob.

OpenStudy (anonymous):

ok thanks